# MDPO: Overcoming the Training-Inference Divide of Masked Diffusion Language Models

### v1

@@ Abstract
Diffusion language models, as a promising alternative to traditional autoregressive (AR) models, enable faster generation and richer conditioning on bidirectional context. However, they suffer from a key discrepancy between training and inference: during inference, MDLMs progressively reveal the structure of the generated sequence by producing fewer and fewer masked tokens, whereas this structure is ignored in training as tokens are masked at random. Although this discrepancy between training and inference can lead to suboptimal performance, it has been largely overlooked by previous works, leaving closing this gap between the two stages an open problem. To address this, we frame the problem of learning effective denoising trajectories as a sequential decision-making problem and use the resulting framework to apply reinforcement learning. We propose a novel Masked Diffusion Policy Optimization (MDPO) to exploit the Markov property diffusion possesses and explicitly train the model under the same progressive refining schedule used at inference. MDPO matches the performance of the previous state-of-the-art (SOTA) method with 60× fewer gradient updates, while achieving average improvements of 9.6% on MATH500 and 54.2% on Countdown over SOTA when trained within the same number of weight updates. Additionally, we improve the remasking strategy of MDLMs as a plug-in inference replacement to overcome the limitation that the model cannot refine tokens flexibly. This simple yet effective training-free strategy, what we refer to as RCR, consistently improves performance and yields additional gains when combined with MDPO. Our findings establish great potential for investigating the discrepancy between pre-training and inference of MDLMs. Code: https://github.com/autonomousvision/mdpo. Project Page: https://cli212.github.io/MDPO/.
@@ 1 Introduction
Remarkable advances in language modeling have largely been driven by the autoregressive (AR) paradigm (Hochreiter & Schmidhuber, 1997; Vaswani et al., 2017), where tokens are generated in a causal fashion. An alternative line of work, Diffusion Language Models (DLMs) generates by iteratively denoising, enabling parallel prediction, bidirectional conditioning and faster generation speed. Recently, DLMs such as Mercury Coder (Khanna et al., 2025), Google Gemini Diffusion (Google DeepMind, 2025), and Seed Diffusion (Song et al., 2025) have emerged as promising alternatives that excel at external math and coding benchmarks. Built upon theoretical insights and practical optimizations in the discrete diffusion framework (Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2025), Masked Diffusion Language Models (MDLMs) leads the current development of DLMs. Diffusion gradually adds noise to data in the forward process, and the model learns to reverse this corruption in the backward process. In the forward process of MDLMs, tokens are randomly chosen to either remain stationary or transition to a predefined absorbing state, such as a special masking token, motivated by the success of BERT (Devlin et al., 2019). Then in the backward process, the model learns to invert the noise by predicting all masked tokens conditioned on partially observed (unmasked) context. For generating a sequence of tokens, MDLMs alternates between predicting masked tokens and selectively remasking a fraction of the predictions based on heuristics and uncalibrated token-wise scores.. Representative works as LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025) demonstrate competitive performance to their similarly sized AR counterparts.
@@
Despite promising empirical results, we identify two fundamental, largely overlooked problems of MDLMs that hinder effective denoising trajectories. (1) The training–inference divide: at inference time, MDLMs follow a model-dependent, confidence-guided unmasking schedule that progressively reveals structure of the generated sequence. By contrast, training masks tokens randomly at sampled diffusion steps, ignoring this progressive schedule. This mismatch prevents the model from learning effective denoising trajectories and often causes Answer Backslide: correct intermediate solutions that are ‘refined’ into wrong final outputs. (2) The common remasking approaches freeze tokens once predicted, making it impossible to revise early low-confidence predictions later. We address these problems with complementary methods in this work. MDPO casts denoising as a multi-step decision-making process, explicitly training the model under the same progressive, model-dependent unmasking schedule used at inference. By using reinforcement learning with intermediate rewards to optimize these trajectories, MDPO aligns training with actual inference dynamics, closing the training–inference divide. Running Confidence Remasking (RCR) is a training-free decoding strategy that tracks per-position confidence over time, enabling flexible remasking and revision of low-confidence tokens at any step. In summary, our contributions are three-fold:
@@
  • Answer Backslide: We observe a novel and previously unknown phenomenon in MDLMs: models occasionally produce correct answers at intermediate steps but ‘refine’ them into incorrect results. We refer to this phenomenon as Answer Backslide, which inspires us to supervise models not only on the final results but also intermediate denoising steps.
@@ 2 Related Work
2.1 Diffusion Language Models
Diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have become the de facto standard for generative modeling in continuous domains, including images, videos, 3D shapes, and robotic trajectories. Building on early efforts in discrete diffusion (Austin et al., 2021; Hoogeboom et al., 2021), several works have adapted diffusion models for text generation. Diffusion-LM (Li et al., 2022) and Self-Conditioned Embedding Diffusion (Strudel et al., 2022) have pioneered embedding-based diffusion approaches, achieving comparable performance to AR methods. Liu et al. (2025) introduces a novel framework that separates the generation process into two models: a planner and a denoiser. Meanwhile, masked diffusion has been established as a special case of discrete diffusion processes (Austin et al., 2021; Sahoo et al., 2024; Ou et al., 2025), with recent research significantly scaling their capabilities. Notably, DiffuLLaMA (Gong et al., 2025) initializes masked diffusion language models using pretrained LLaMA weights. Block Diffusion (Arriola et al., 2025) models sequences in autoregressive blocks with intra-block diffusion, enabling variable-length generation and faster inference via KV caching. LLaDA (Nie et al., 2025) and Dream (Ye et al., 2025) scale MDLMs by pre-training from scratch to the billion-parameter regime, achieving performance comparable to similarly sized AR LLMs.
Beyond scaling, diffusion models have also shown promise in reasoning tasks. Ye et al. (2024) demonstrates that DLMs are effective in producing chain-of-thought reasoning and addressing complex reasoning tasks, even with smaller-scale models. These findings highlight diffusion models as a viable alternative to traditional AR methods. Kim et al. (2025); Wang et al. (2025) show that though MDLMs face inherently harder training challenges due to solving many computationally intractable infilling tasks, using adaptive decoding strategies dramatically overcomes these limitations. Our proposed Running Confidence Remasking builds upon these and shows that enabling the model to iteratively revise early-predicted ‘noisy’ tokens substantially enhances downstream performance.
2.2 Training Diffusion Models with Reinforcement Learning
Reinforcement learning has been primarily explored in training image diffusion models (Black et al., 2024; Fan et al., 2023) on non-differentiable downstream objectives such as human-perceived image quality or drug effectiveness. RL has emerged as a promising tool to optimize such sparse and task-specific reward signals (Yu et al., 2025; Shao et al., 2024; Luo et al., 2025b; a; Hu et al., 2025; Zeng et al., 2025). In language modeling, reinforcement learning has been extensively applied to AR LLMs, where post-training with reward-based fine-tuning significantly boosts reasoning capabilities (Guo et al., 2025; Luo et al., 2025b; a; Zeng et al., 2025; Hu et al., 2025). These RL algorithms are tailored for sequential token-by-token generation, a key property of AR models. However, MDLMs follow a fundamentally different paradigm: they are trained to predict complete text generations given randomly corrupted text, and iteratively refine the generations with fewer and fewer masked tokens at inference time. This prevents RL algorithms designed for AR models from being directly applied to DLMs. Zhao et al. (2025) made the first effort in this direction by adapting pretrained MDLMs to RL using Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a policy optimization algorithm originally designed for AR models. In contrast, our work introduces MDPO, a policy optimization method designed for MDLMs to learn effective denoising trajectories by optimizing the iterative denoising process as a multi-step decision-making problem. To our knowledge, MDPO is the first RL method specifically designed to optimize the denoising trajectory of MDLMs.
@@ 3 Methodology
3.1 Masked Language Diffusion Models
We first provide the formulation of LLaDA (Nie et al., 2025), a pre-trained MDLM that serves as the foundation of our work. While specific architectural or methodological variations may exist among other MDLMs, the methods proposed in this paper are broadly applicable to the majority of existing MDLM frameworks. LLaDA defines a model distribution $p_{\theta}(x_{0})$ of the clean data $x_{0}$ through a discrete diffusion process (Ou et al., 2025; Austin et al., 2021). The forward process gradually masks tokens until the sequence is fully masked. The reverse process recovers $p_{\theta}(x_{0})$ by iteratively predicting masked tokens.
Training: The training objective optimizes a mask predictor $p_{\theta}\left(x_{0}\mid\bar{x}_{\tau}\right)$ that takes partially masked data $\bar{x}_{\tau}$ [^1] We use the bar notation (e.g. $\bar{x}$) to denote masked and $x$ for unmasked sequences throughout the paper. as input and predicts all masked tokens simultaneously using a cross-entropy loss:
@@ 3.1 Masked Language Diffusion Models
    Given a masked sequence $\bar{x}_{t}$, draw all masked tokens by sampling from $x_{t-1}\sim p_{\theta}\big{(}\cdot\mid\bar{x}_{t}\big{)}$ to produce a fully unmasked sequence $x_{t-1}$.
@@
    Remask $x_{t-1}$ using a confidence-based remasking function $\gamma$ yielding a partially masked sequence $\bar{x}_{t-1}$. More specifically, with decreasing $t$, LLaDA remasks a decreasing number of tokens based on the current model’s confidence score $1-p_{\theta}(x_{t-1}^{i}\mid\bar{x}_{t})$
LLaDA continues this process until time step $t=0$. The confidence-based remasking strategy (2) reveals more and more structure for $p_{\theta}(x_{t-1}|\bar{x}_{t})$ to condition on over time. The number of denoising steps $T$ is a hyperparameter providing a trade-off between efficiency and quality (Nie et al., 2025). We refer the reader to Section 3.4 for details about the implementation of the remasking function $\gamma$.
@@ 3.2 The Training-Inference Divide
Table 1: Accuracy of LLaDA on two verifiable tasks and the ratio of at least one out of 256 intermediate steps is correct (w/ inter.ans).
The above formulation of LLaDA reveals an important discrepancy between training and inference: specifically, during training, the model is optimized to predict all masked tokens at randomly chosen masking ratios $\tau$. In contrast, inference involves multiple iterative denoising steps informed by the model’s confidence score, forming a trajectory with fewer and fewer masked tokens to be predicted, and more and more structure being revealed. However, this structure is ignored during training, where at each iteration tokens are masked at random. This leads to a training-inference divide which limits the performance of the model as it is unable to learn effective diffusion trajectories.
We observe symptoms of this behavior by inspecting denoising trajectories of pre-trained MDLMs. Table 1 shows the accuracy of LLaDA on two reasoning tasks and the additional accuracy if we count intermediate-step correct answers that are ‘refined’ into wrong final answers. We refer to this phenomenon as Answer Backslide and do detailed studies in Section 4.3.
@@
A key advantage of masked language diffusion models over traditional auto-regressive models is that they yield complete text generations at every inference step. This allows for evaluating the quality of intermediate generations using an appropriately chosen reward model. In this paper, we propose to exploit this property to effectively fine-tune diffusion policies using deep reinforcement learning. To focus our analysis, we investigate reasoning tasks with verifiable answers, such as math problems. However, we remark that other problems without verifiable answers can also be addressed within our framework using modern validation concepts such as LLM-as-a-judge (Li et al., 2024).
@@ 3.3 Masked Diffusion Policy Optimization (MDPO)
The expected return of our Masked Diffusion Policy Optimization (MDPO) model is:
@@
where the expectation is computed over denoising trajectories $\{x_{T},\dots,x_{0}\}$ generated by the current denoising policy $p_{\theta}$. We use the same masking function $\gamma(x_{t},t)$ as used during inference, which takes an unmasked sequence $x_{t}$ as input and converts it into a (partially) masked sequence $\bar{x}_{t}$ depending on the current time step $t$. The smaller the timestep, the fewer tokens are masked (see Section 3.4 for details). The indicator function aggregates the return only at locations where the input is masked. The denoising policy models the probability distribution of the denoised token $x^{i}_{t-1}$ at location $i$ given the masked input at the previous time step $\bar{x}_{t}$. Again, we drop the prompt here for notational clarity.
@@
Here, $p_{\text{ref}}$ is the reference model which is usually the initial model before training, $\mathbb{D}_{KL}\left[p_{\theta}||p_{\text{ref}}\right]$ is the KL divergence between the trained model $p_{\theta}$ and the reference model. $Z_{t,k}^{i}$ is the clipped and weighted advantage (Schulman et al., 2017)
@@
where $\epsilon$ and $\beta$ are hyper-parameters. $A_{t,k}$ is the advantage of $x_{t}$ in the $k$’th group calculated based on relative rewards inside each group:
@@
In GRPO, group-relative advantage estimation optimizes the memory usage of training LLMs with RL by dropping the jointly trained value model that is originally used to predict the advantages. Beyond this, another intuitive reason to use group-relative advantage estimation in the diffusion setting is to incentivize better generation with fewer denoising steps. Specifically, by comparing to a group of trajectories, steps reaching the final reward more rapidly are assigned higher advantages. In addition, Eq. (6) encourages each denoising step to (i) make an immediate improvement over the previous, noisier prediction and (ii) steer the trajectory toward cleaner states that score well overall.
@@ 3.4 Remasking
As introduced in Section 3.1, inference in LLaDA alternates between predicting all masked tokens in $\bar{x}_{t}$ and selectively re-masking a fraction of the prediction $x_{t-1}$ via $\bar{x}_{t-1}=\gamma(x_{t-1},t)$ for iterative refinement. We first discuss random and confidence-based re-masking as introduced in LLaDA. Next, we propose a simple yet effective improvement to address a key limitation of confidence-based re-masking, which consistently leads to better performance for both LLaDA pre-trained as well as MDPO fine-tuned models. Formally, the remasking function $\gamma$ assigns the mask symbol $\mathcal{M}$ to the tokens with the $n_{t}$’th highest masking scores:
@@
where $n_{t}=\lfloor\phi(t,T)L\rceil$ determines the number of tokens to be masked depending on the time step $t$ and a scheduling function $\phi$. $n_{t}$ decreases as $t$ increases (more tokens are masked for larger $t$). $m_{t}^{i}$ is a generic definition of the masking score of token $i$ at step $t$ and remasking strategies vary in how they compute $m_{t}^{i}$. LLaDA applies a linear function for $\phi$ so that an equal amount of $\lfloor\frac{L}{T}\rceil$ tokens are not remasked every step, i.e., $n_{t}=\lfloor\frac{T-t}{T}L\rceil$. We investigate various schedules in Appendix A.2.
@@
Low-Confidence Remasking (LCR): Inspired by the annealing tricks of sampling in AR LLMs (Holtzman et al., 2020), LLaDA proposes to utilize the model’s prediction confidence as the masking scores:
@@
Figure 2: Comparison between Lowest Confidence Remasking (LCR) and our proposed Running Confidence Remasking (RCR) during iterative denoising. For each step, we show tokens that are not remasked. LCR freezes low-confidence tokens once unmasked, preventing further refinement, which potentially accumulates early-stage noise. For example, the token ‘problem’ predicted and frozen by LCR at step 1 is wrong but maintained until the end of the denoising, which leads to the final wrong answer. Whereas RCR tracks the running maximum confidence for each position, allowing persistently low-confidence tokens to be refined in later steps, leading to higher-quality completions.

### v2

@@ Abstract
Diffusion language models, as a promising alternative to traditional autoregressive (AR) models, enable faster generation and richer conditioning on bidirectional context. However, they suffer from a key discrepancy between training and inference: during inference, MDLMs progressively reveal the structure of the generated sequence by producing fewer and fewer masked tokens, whereas this structure is ignored in training as tokens are masked at random. Although this discrepancy between training and inference can lead to suboptimal performance, it has been largely overlooked by previous works, leaving closing this gap between the two stages an open problem. To address this, we frame the problem of learning effective denoising trajectories as a sequential decision-making problem and use the resulting framework to apply reinforcement learning. We propose a novel Masked Diffusion Policy Optimization (MDPO) to exploit the Markov property diffusion possesses and explicitly train the model under the same progressive refining schedule used at inference. MDPO matches the performance of the previous state-of-the-art (SOTA) method with 60× fewer gradient updates, while achieving average improvements of 9.6% on MATH500 and 54.2% on Countdown over SOTA when trained within the same number of weight updates. Additionally, we improve the remasking strategy of MDLMs as a plug-in inference replacement to overcome the limitation that the model cannot refine tokens flexibly. This training-free method, termed Running Confidence Remasking (RCR), consistently enhances performance and provides further improvements when used with MDPO. Our findings establish great potential for investigating the discrepancy between pre-training and inference of MDLMs. Code: https://github.com/autonomousvision/mdpo. Project Page: https://cli212.github.io/MDPO/.
@@ 1 Introduction
Remarkable advances in language modeling have largely been driven by the autoregressive (AR) paradigm (Hochreiter & Schmidhuber, 1997; Vaswani et al., 2017), where tokens are generated causally. Diffusion Language Models (DLMs), as an alternative line of work, generate by iterative denoising to enable parallel prediction, bidirectional conditioning, and faster generation speed. Recently, DLMs such as Mercury Coder (Khanna et al., 2025), Google Gemini Diffusion (Google DeepMind, 2025), and Seed Diffusion (Song et al., 2025) have emerged as promising alternatives that excel at external math and coding benchmarks. Built upon theoretical insights and practical optimizations in the discrete diffusion framework (Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2025), Masked Diffusion Language Models (MDLMs) lead the current development of DLMs. In the forward process of MDLMs, tokens are randomly chosen to either remain stationary or transition to a predefined absorbing state, such as a special masking token, motivated by the success of BERT (Devlin et al., 2019). Then, in the backward process, the model learns to invert the noise by predicting all masked tokens conditioned on partially observed (unmasked) context. For generating a sequence of tokens, current MDLMs alternate between predicting masked tokens and selectively remasking a fraction of the predictions based on heuristics and uncalibrated token-wise scores. Representative works such as LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025) demonstrate competitive performance to their similarly sized AR counterparts.
@@
Despite promising empirical results, we identify two fundamental, largely overlooked problems of MDLMs that hinder effective denoising trajectories. (1) The training–inference divide: at inference time, MDLMs follow a model-dependent, confidence-guided remasking schedule that progressively reveals the structure of the generated sequence. By contrast, tokens are randomly masked at sampled diffusion steps in training, ignoring the progressive schedule. This mismatch prevents the model from learning effective denoising trajectories and often causes over-denoising: correct intermediate solutions are ‘refined’ into wrong final solutions (see Fig. 3 for statistical analysis). (2) The common remasking approaches freeze tokens once predicted, making it impossible to revise early low-confidence predictions at later denoising steps. We address these problems with complementary methods in this work. Masked Diffusion Policy Optimization (MDPO) casts denoising as a multi-step decision-making process, explicitly training the model under the same progressive, model-dependent remasking schedule used at inference. By using reinforcement learning with intermediate rewards to optimize these trajectories, MDPO aligns training with actual inference dynamics, closing the training–inference divide. Running Confidence Remasking (RCR) is a training-free decoding strategy that tracks per-position confidence over time, enabling flexible remasking and revision of low-confidence tokens at any step. In summary, our contributions are threefold:
@@
  • Over-denoising: We observe a novel and previously unknown phenomenon in MDLMs: models occasionally produce correct answers at intermediate steps but ‘refine’ them into incorrect results. We refer to this phenomenon as over-denoising, which inspires us to supervise models not only on the final results but also intermediate denoising steps.
@@ 2 Methodology
2.1 Preliminary: Masked Diffusion Language Models
We introduce our methods based on the formulation of LLaDA (Nie et al., 2025), a pre-trained MDLM that models clean data $x_{0}$ via a discrete diffusion process (Ou et al., 2025; Austin et al., 2021), where tokens are progressively masked in the forward process and then recovered by iterative prediction in the reverse process. While specific methodological variations may exist, the methods proposed in this paper are broadly applicable to the majority of existing MDLM frameworks.
Training: The training objective optimizes a mask predictor $p_{\theta}\left(x_{0}\mid\bar{x}_{\tau}\right)$ that takes partially masked data $\bar{x}_{\tau}$ [^1] We use the bar notation $\bar{x}$ to denote (partially) masked and $x$ for unmasked sequences throughout the paper. as input and predicts all masked tokens simultaneously using a cross-entropy loss:
@@ 2.1 Preliminary: Masked Diffusion Language Models
    Given a masked sequence $\bar{x}_{t}$, draw all masked tokens by sampling $x_{t-1}\sim p_{\theta}\big(\cdot\mid\bar{x}_{t}\big)$ to produce a fully unmasked sequence $x_{t-1}$.
@@
    Remask $x_{t-1}$ using a confidence-based remasking function $\gamma$ yielding a partially masked sequence $\bar{x}_{t-1}$. More specifically, with decreasing $t$, LLaDA remasks a decreasing number of tokens based on the current model’s confidence score $p_{\theta}(x_{t-1}^{i}\mid\bar{x}_{t})$
LLaDA continues this process until $x_{0}$ is sampled. The confidence-based remasking strategy in the second step above reveals more and more structure for $p_{\theta}(\cdot\mid\bar{x}_{t})$ in the first step to condition on over time. $T$ is a hyperparameter that trades off computation and quality (Nie et al., 2025) in practice. We refer readers to Section 2.4 for details about the implementation of the remasking function $\gamma$.
@@ 2.2 The Training-Inference Divide
Table 1: Accuracy of LLaDA on verifiable tasks and the ratio that at least one of 256 intermediate steps is correct (w/ inter. ans).
The above formulation of LLaDA reveals an important discrepancy between training and inference: During training, the model is optimized to predict all masked tokens at randomly sampled masking ratios $\tau$. In contrast, inference involves multiple iterative denoising steps informed by the model’s confidence score, forming a trajectory with fewer and fewer masked tokens to be predicted, and more and more structure being revealed. However, this structure is ignored during training where tokens are masked at random at each iteration. This training-inference divide limits the performance of the model as the model is unable to learn effective denoising trajectories.
We observe symptoms of this divide by inspecting denoising trajectories of pre-trained MDLMs. Table 1 shows the accuracy of LLaDA on two reasoning tasks and the additional accuracy if we count intermediate-step correct answers that are ‘refined’ into wrong final answers. We term this phenomenon as over-denoising and conduct detailed studies in Section 3.3.
@@
A key advantage of masked diffusion language models over traditional auto-regressive models is that they yield complete text generations at every inference step. This allows for evaluating the quality of intermediate generations using an appropriately chosen reward model. In this paper, we propose to exploit this property to effectively fine-tune diffusion policies using deep reinforcement learning. To focus our analysis, we investigate reasoning tasks with verifiable answers, such as math problems. However, we remark that other problems without verifiable answers can also be addressed within our framework using modern validation concepts such as LLM-as-a-judge (Li et al., 2024).
@@ 2.3 Masked Diffusion Policy Optimization (MDPO)
Based on this, we derive the expected return of our Masked Diffusion Policy Optimization:
@@
where the expectation is computed over denoising trajectories $\{x_{T},\dots,x_{0}\}$ generated by the current denoising policy $p_{\theta}$. We use the same masking function $\gamma(x_{t},t)$ as used during inference, which takes an unmasked sequence $x_{t}$ as input and converts it into a (partially) masked sequence $\bar{x}_{t}$ conditioned on the current time step $t$. The smaller the timestep, the fewer tokens are masked (see Section 2.4 for details). The indicator function aggregates the return only at locations where the input is masked. The denoising policy models the probability of the denoised token $x^{i}_{t-1}$ at location $i$ given the masked input at the previous time step $\bar{x}_{t}$. Again, we drop the prompt here for notational clarity.
@@
Here, $p_{\text{ref}}$ is the reference policy which is usually the initial model before training. $\mathbb{D}_{KL}\left[p_{\theta}||p_{\text{ref}}\right]$ is the KL divergence between the trained policy $p_{\theta}$ and the reference policy to avoid large deviation. We define the clipped and weighted advantage (Schulman et al., 2017) as
@@
where $\epsilon$ and $\beta$ are hyper-parameters. $A_{t,k}$ is the advantage of $x_{t,k}$, the $k$’th rollout in the group, calculated based on relative rewards inside the group:
@@
The original group-relative advantage estimation optimizes the memory usage of training LLMs with RL by dropping the jointly trained value model that is used to predict the advantages. Beyond this, another intuitive reason to use group-relative advantage estimation in the diffusion setting is to incentivize better generation with fewer denoising steps. Specifically, by comparing to a group of trajectories, steps reaching the final reward more rapidly are assigned higher advantages. In addition, Eq. (6) encourages each denoising step to (i) make an immediate improvement over the previous, noisier prediction and (ii) steer the trajectory toward cleaner states that score well overall.
@@ 2.4 Remasking
As introduced in Section 2.1, inference in LLaDA alternates between predicting all masked tokens in $\bar{x}_{t}$ and selectively remasking a fraction of the prediction $x_{t-1}$ via $\bar{x}_{t-1}=\gamma(x_{t-1},t)$ for iterative refinement. We first discuss random and confidence-based remasking as introduced in LLaDA. Next, we propose a simple yet effective improvement to address a key limitation of confidence-based remasking, which consistently leads to better performance for both LLaDA pre-trained as well as MDPO fine-tuned models. Formally, the remasking function $\gamma$ assigns the mask symbol $\mathcal{M}$ to the $n_{t}$ tokens with the highest masking scores:
@@
where $n_{t}=\lfloor\phi(t,T)L\rceil$ determines the number of tokens to be masked depending on the time step $t$ and a scheduling function $\phi$. $n_{t}$ decreases as $t$ increases (more tokens are masked for larger $t$). $m_{t}^{i}$ is a generic definition of the masking score of token $i$ at step $t$ and remasking strategies vary in how they compute $m_{t}^{i}$. LLaDA applies a linear function for $\phi$ so that an equal amount of $\lfloor\frac{L}{T}\rceil$ tokens are not remasked every step, i.e., $n_{t}=\lfloor\frac{T-t}{T}L\rceil$. We investigate various schedules in Appendix B.2.
@@
Low-Confidence Remasking (LCR): Inspired by the annealing tricks of sampling in AR LLMs (Holtzman et al., 2020), LLaDA proposes to utilize the complement of the predicted confidence of the model as the masking scores:
@@
Figure 2: Comparison between Low-Confidence Remasking (LCR) and our proposed Running Confidence Remasking (RCR) during iterative denoising. For each step, we show tokens that are not remasked. LCR freezes low-confidence tokens once unmasked, preventing further refinement, which potentially accumulates early-stage noise. For example, the token ‘problem’ predicted and frozen by LCR at step 1 is wrong but maintained until the end of the denoising, which leads to the final wrong answer. Whereas RCR tracks the running maximum confidence for each position, allowing persistently low-confidence tokens to be refined in later steps, leading to higher-quality completions.